OpenStudy (anonymous):

Ned ordered 3/4 of a ton of gravel. He received 2/3 of a ton of gravel. what fraction of his order did he receive? @Mathematics

OpenStudy (anonymous):

my question asks me to draw a picture to represent this... that is where I am struggling  this is for a math for teachers class and I need to represent the whole

OpenStudy (stormfire1):

Oops...I read that wrong. He ordered 3/4 and got 2/3. So it's really \[\frac{\frac{2}{3}}{\frac{3}{4}}\]Then flip things around to get:\[\frac{2}{3}*\frac{4}{3}=\frac{8}{9}\]

OpenStudy (anonymous):

that is what I thought but I was just having trouble drawing it and having it make sense!

OpenStudy (mkapad01):

\[3/4\div2/3\]

OpenStudy (stormfire1):

You can also look at it this way: he ordered .75 and got .67: .67/.75 = ~.89 on a calculator...but the exact answer is 8/9

OpenStudy (mkapad01):